A 1.55-mol sample of nitrogen gas is maintained in a 0.730-L container at 292 K. Calculate the pressure of the gas using both the ideal gas law and the van der Waals equation (van der Waals constants for N2 are a = 1.39 L2atm/mol2 and b = 3.91×10-2 L/mol).

Pideal gas equation = ______ atm

Pvan der Waals =_____ atm
Answer #1

1)

Given:

V = 0.73 L

n = 1.55 mol

T = 292.0 K

use:

P * V = n*R*T

P * 0.73 L = 1.55 mol* 0.08206 atm.L/mol.K * 292 K

P = 50.9 atm

Answer: 50.9 atm

2)

Given:

V = 0.73 L

n = 1.55 mol

R = 0.08206 atm.L/mol.K

T = 292.0 K

a = 1.39 atm.L^2/mol^2

b = 0.0391 L/mol

use:

(P+an^2/V^2)*(V-nb) = n*R*T

(P + 1.39*1.55^2/0.73^2)*(0.73-1.55*0.0391) = 1.55*0.08206*292.0

(P + 6.2666)*(0.6694) = 37.1404

P + 6.2666 = 55.4835

P = 49.2 atm

Answer: 49.2 atm